OpenStudy (anonymous):

A set contains all possible five digit positive integers that can be made using each of the digits 1,3,5,7, and 9 once. Ex of possible integers: 13579; 35791;etc. What is the median of that set? (not the example, but all of it)

OpenStudy (anonymous):

Pleeeease help, someone????

OpenStudy (anonymous):

I would start by subtracting the smallest possible number from the largest and then dividing by 2. But to premise that, is the smallest number 11111 or is it 13579?

OpenStudy (jim766):

can only use each digit once

OpenStudy (jim766):

13579 would be first 97531 would be last

OpenStudy (jim766):

wouldn't the number of possible numbers using 5 numbers 5*4*3*2*1 ?

OpenStudy (anonymous):

Drat. My brilliant plan has resulted in 41976 I think I may have bitten off more than I can chew here... :-(

OpenStudy (anonymous):

I know, this problem is hard and I'm trying my best to solve it and ask others too:)

jimthompson5910 (jim_thompson5910):

well we know this for sure there are 5*4*3*2*1 = 120 different ways to arrange the 5 numbers

jimthompson5910 (jim_thompson5910):

the median will be in between locations 60 and 61 (the smallest number 13579 is at location 1 and the largest number 97531 is at location 120)

jimthompson5910 (jim_thompson5910):

actually it will be the average of the values at locations 60 and 61

OpenStudy (anonymous):

Thank you:)

jimthompson5910 (jim_thompson5910):

the first 24 values start with a 1 (since 4! = 4*3*2*1 = 24) the next 24 values start with a 3 so that's 24+24 = 48 values so far that are below the median

jimthompson5910 (jim_thompson5910):

so the median must start with a 5 because 48+24 = 72 telling us that we've overshot locations 60 and 61

OpenStudy (anonymous):

That makes sense now:) Thanks!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

@jim_thompson5910 You have a knack for bringing highly illuminating forehead slappers to the table! :-) This is really interesting, and after studying your claims, I'll try reiterating them (verbosely) in a way I hope will help others to understand. I only hope this won't make things MORE confusing to anyone. As you said, There are 5*4*3*2 = 120 ways of arranging the digits: 1,3,5,7,9 If you divide these 120 numbers into groups based on their first digit, you'll find there are 120/5 = 24 numbers in each group starting with a 1,3,5,7 or 9. So since there are 24 numbers in the 1nnnn group and 24 numbers in the 3nnnn group, we know that the 60th number must be in the 5nnnn group. Now this is where Jim left off, but it turns out we can also use Jim's logic to continue on to find the second digit in the 5nnnn group: Since we know that the 5-digit number we want starts with (and uses up) the digit "5", then we also know there are 4*3*2 = 24 ways of arranging the remaining digits: 1,3,7,9 And just as before, if you divide these 24 numbers into groups based on their first digit, you'll find there are 24/4 = 6 numbers in each group starting with a 1,3,7 or 9. So since there are 6 numbers in the 51nnn group and 6 numbers in the 53nnn group, and that 24+24+6+6=60, then we know that the 60th number must be the last number in the 53nnn group, which would make it 53971. But since the median is the "middle" value in a sequence, the "middle" of a list of 120 numbers lies between the 60th and the 61st number. This means the median must be artificially created by averaging the 60th and 61st number. So since the 60th number was the LAST number in the 53nnn group, then the 61st number must be the FIRST number in the following group which is 57nnn, and the first number in that group would be 57139. 1 13579 2 13597 3 13759 .. ... 56 ... 57 53719 58 53791 59 53917 60 53971 - last number in the 53nnn group 60.5 - median position in a list of 120 numbers 61 57139 - first number in the 57nnn group 62 57193 63 57319 64 57391 65 ... .. ... 118 97351 119 97513 120 97531 So bottom line here is that I figure the median would be (53971+57139)/2 Numbers check anyone? :-)

jimthompson5910 (jim_thompson5910):

that's clever, didn't think to extend it but that makes sense now that I think about it and yes, the median is 55,555

OpenStudy (anonymous):

YAY!!, as they say... :-) And thank you!

jimthompson5910 (jim_thompson5910):

you're welcome